# Nested ensemble Kalman filter for static parameter inference in nonlinear state-space models

### v1

@@ 5.3 Two-node epidemic model
We consider here a two-node compartmental epidemic model to characterise the spread of oak processionary moth (OPM) in two London parks (Bushy Park and Richmond Park). Similar to wadkin23, we specify a stochastic susceptible-infected-removed (SIR) model (e.g. AnBr00) for each node (park), with common removal rate $\gamma$ and infestation ‘pressure’ from the neighbouring node, described by the parameters $\alpha_{ij}$, where $\alpha_{12}$ is the pressure applied by node 1 (Bushy) on node 2 (Richmond), and $\alpha_{21}$ is the pressure from node 2 on node 1. Unlike wadkin23, we allow time-varying infestation rates $\beta_{1,t}$ and $\beta_{2,t}$ at each node, described by two independent Brownian motion processes, scaled by a common parameter $\sigma_{\beta}$. The model takes the form of (25), with $X_{t}=(S_{1,t},I_{1,t},S_{2,t},I_{2,t},\log\beta_{1,t},\log\beta_{2,t})^{\prime}$ and $\theta=(\gamma,\alpha_{1,2},\alpha_{2,1},\sigma_{\beta})^{\prime}$. The drift and diffusion functions are given by
@@
The data (see Figure 5) consist of (noisy) observations of removal prevalence $(R_{1,t},R_{2,t})^{\prime}$, that is, the cumulative numbers of tree locations where nest removal has taken place, observed each year over the period 2013-2024. We assume a fixed number of tree locations ($\text{N}_{j}$) at each node $j$ so that $S_{j,t}+I_{j,t}+R_{j,t}=\text{N}_{j}$, $j=1,2$, and observation of removals is equivalent to observation of the sums $S_{1,t}+I_{1,t}$ and $S_{2,t}+I_{2,t}$. Following wadkin23, we specify an observation model with density
@@
We adopted an independent prior specification with $\text{Gamma}(2,2)$, for all parameter components except $\sigma_{\beta}$, for which we ascribed a $\text{Gamma}(2,10)$ distribution, reflecting our prior belief that the intrinsic stochasticity in the infestation process is small, relative to the observation noise. We ran NEnKF and RB-SMC2, both of which use Algorithm 3, which requires a linear Gaussian approximation of (28). For NEnF, we used $R=\sigma^{2}\text{diag}\{H\hat{\mu}_{t|t-1}\}$; for RB-SMC2, this quantity was further scaled by a factor of 2. We performed a single run of these schemes with $M=20,000$ particles and $N=20$ initial ensemble members. An initial condition of $x_{0}=(4631,240,37413,1400,-10,-10.5)^{\prime}$ was assumed to be fixed and known for each run. We followed Section 5.2 by sampling $p_{t}^{(\theta)}(x_{t}|x_{t-1})$ through recursive draws of the Euler-Maruyama approximation (26) with $\Delta t=0.1$. The conditions for triggering a resample-move step and increasing $N$ were as Section 5.1. Both schemes used a delayed acceptance kernel in the move step (with $k=3$), which gave a decrease in CPU time of approximately 15% for NEnKF and 25% for RB-SMC2, over use of a standard (pseudo-marginal) MH kernel.
@@ 5.4 Lorenz 96
with addition and subtraction of component indices interpreted modulo $d$. The diffusion matrix is a diagonal matrix such that $\Sigma=\theta_{4}^{2}I_{d\times d}$. This SDE can be seen as a stochastic representation of the Lorenz 96 dynamical system (Lorenz1996), and is often used for benchmarking inference schemes in high-dimensional settings. Following drovandi22, we set $x_{0}=(0,\ldots,0)^{\prime}$ and $\theta=(1,1,8,\sqrt{10})^{\prime}$. Two synthetic data sets, $\mathcal{D}_{1}$ (with $d=5$) and $\mathcal{D}_{2}$ (with $d=10$), were generated via recursive sampling of the Euler-Maruyama approximation of (29) over $[0,30]$ with $\Delta t=5\times 10^{-3}$. The resulting paths were thinned by a factor of 40 and corrupted with additive Gaussian noise, with zero mean and variance $R=5^{2}I_{d\times d}$, giving, for each of $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$, 30 observations with an inter-observation time of $0.2$ time units.
@@ 6 Discussion
Bayesian filtering for state-space models in the presence of unknown static parameters is a challenging problem. Sequential Monte Carlo algorithms must overcome particle degeneracy; one such scheme that deals with static parameters in a principled way is the SMC2 scheme of Chopin2013. Here, parameter particles are weighted by an estimate of the current observed-data likelihood and, subject to some degeneracy criterion, mutated via a (pseudo-marginal) Metropolis-Hastings kernel which has the marginal parameter posterior as its invariant distrubution. A particle filter over the latent state process is used to estimate the required observed-data likelihood contributions. However, use of the particle filter in this way is likely computationally prohibitive in high-dimensional observation settings, due to requiring that the number of state particles be scaled exponentially with dimension. On the other hand, the ensemble Kalman filter (EnKF, evensen1994) shifts particles (known as ensemble members in the EnKF context) via suitable Gaussian approximations, thus avoiding resampling altogether. Consequently, there is now a vast literature demonstrating the empirical performance of the EnKF for high-dimensional state-space models (see e.g. Katzfuss2019, and the references therein).
@@
We have proposed to use the EnKF within the principled framework of SMC2 by weighting parameter particles by the EnKF likelihood and, when necessary, mutating these particles by using a Metropolis-Hastings step that again uses the EnKF likelihood. Extensions to the resulting nested EnKF (NEnKF) algorithm were considered. In particular, the computational cost of the mutation step can be lessened by screening proposed moves using a surrogate, computed locally using the most recent parameter samples and their likelihood evaluations. As with the simplest implementation of SMC2, NEnKF is plug-and-play, since only forward simulation of the latent state process and evaluation of the observation density are required. We found, across four applications of increasing observation dimension, that the NEnKF is able to provide accurate and computationally efficient inferences regarding the static parameters. Since the NEnKF combines a particle filter over parameters with an ensemble Kalman filter over states, we anticipate (and as evidenced by our numerical experiments) greatest utility of the proposed algorithm in settings where the number of observed dynamic components is large compared to the number of parameters to be estimated (e.g. $d_{\theta}<10$ and $d_{y}>10$). The simplest implementation of the NEnKF assumes a linear Gaussian observation model. In scenarios where this assumption is unreasonable, we have proposed to replace the EnKF with a particle filter that uses the EnKF to propagate state particles. Unlike related approaches (e.g. papadakis2010), we also use the EnKF to give a linear Gaussian approximation of the transition density $p_{t}^{(\theta)}(x_{t}|x_{t-1})$, leading to a simple form of weight function that can be readily applied in settings where states are modelled at a finer frequency than observations (as is the case, for example, when working with discretised SDEs).
@@
This work can be extended in a number of ways. For example, in our applications of the delayed acceptance kernel, the number of distinct parameter particles $M_{r}$ used to construct the $k$-$n$ surrogate was typically small (no more than a few thousand particles), justifying an $\mathcal{O}(M_{r})$ list look-up method for finding the $k$-nearest neighbours of each $\theta^{*}$. Using a variation of the KD-tree (e.g. Sherlock2017) requires $\mathcal{O}(d_{\theta}\log M_{r})$ operations and provides a cheaper alternative to the list look-up method, provided $d_{\theta}$ is moderate. It may also be possible to improve computational efficiency of the NEnKF by adaptively choosing between a delayed-acceptance kernel and standard M-H kernel, when the resample-move step is triggered. For example, in the early phase of the NEnKF scheme, calculating an observed-data likelihood estimator is likely to be relatively cheap. Similar approaches have been considered in the context of IBIS (taylor13) and SMC2 (botha23adapt). Finally, using the surrogate in both the weighting and resample-move steps (for a fixed value of $N$) could provide a fixed cost alternative to the NEnKF, to the detriment of inferential accuracy.

### v2

@@ 5.3 Two-node epidemic model
We consider here a two-node compartmental epidemic model to characterise the spread of oak processionary moth (OPM) in two London parks (Bushy Park and Richmond Park). Similar to Wadkin et al. (2023), we specify a stochastic susceptible-infected-removed (SIR) model (e.g. Andersson and Britton, 2000) for each node (park), with common removal rate $\gamma$ and infestation ‘pressure’ from the neighbouring node, described by the parameters $\alpha_{ij}$, where $\alpha_{12}$ is the pressure applied by node 1 (Bushy) on node 2 (Richmond), and $\alpha_{21}$ is the pressure from node 2 on node 1. Unlike Wadkin et al. (2023), we allow time-varying infestation rates $\beta_{1,t}$ and $\beta_{2,t}$ at each node, described by two independent Brownian motion processes, scaled by a common parameter $\sigma_{\beta}$. The model takes the form of (25), with $X_{t}=(S_{1,t},I_{1,t},S_{2,t},I_{2,t},\log\beta_{1,t},\log\beta_{2,t})^{\prime}$ and $\theta=(\gamma,\alpha_{1,2},\alpha_{2,1},\sigma_{\beta})^{\prime}$. The drift and diffusion functions are given by
@@
The data (see Figure 5) consist of (noisy) observations of removal prevalence $(R_{1,t},R_{2,t})^{\prime}$, that is, the cumulative numbers of tree locations where nest removal has taken place, observed each year over the period 2013-2024. We assume a fixed number of tree locations ($\text{N}_{j}$) at each node $j$ so that $S_{j,t}+I_{j,t}+R_{j,t}=\text{N}_{j}$, $j=1,2$, and observation of removals is equivalent to observation of the sums $S_{1,t}+I_{1,t}$ and $S_{2,t}+I_{2,t}$. Following Wadkin et al. (2023), we specify an observation model with density
@@
We adopted an independent prior specification with $\text{Gamma}(2,2)$, for all parameter components except $\sigma_{\beta}$, for which we ascribed a $\text{Gamma}(2,10)$ distribution, reflecting our prior belief that the intrinsic stochasticity in the infestation process is small, relative to the observation noise. We ran NEnKF and RB-SMC2, which use Algorithm 3 and 7 respectively, and require a linear Gaussian approximation of (28). For NEnKF, we used $\tilde{R}=\sigma^{2}\text{diag}\{H\hat{\mu}_{t|t-1}\}$; for RB-SMC2, this quantity was further scaled by a factor of 2. We performed a single run of these schemes with $M=20,000$ particles and $N=20$ initial ensemble members. An initial condition of $x_{0}=(4631,240,37413,1400,-10,-10.5)^{\prime}$ was assumed to be fixed and known for each run. We followed Section 5.2 by sampling $p_{t}^{(\theta)}(x_{t}|x_{t-1})$ through recursive draws of the Euler-Maruyama approximation (26) with $\Delta t=0.1$. The conditions for triggering a resample-move step and increasing $N$ were as Section 5.1. Both schemes used a delayed acceptance kernel in the move step (with $k=3$), which gave a decrease in CPU time of approximately 15% for NEnKF and 25% for RB-SMC2, over use of a standard (pseudo-marginal) MH kernel.
@@ 5.4 Lorenz 96
with addition and subtraction of component indices interpreted modulo $d$. The diffusion matrix is a diagonal matrix such that $\Sigma=\theta_{4}^{2}I_{d\times d}$. This SDE can be seen as a stochastic representation of the Lorenz 96 dynamical system (Lorenz, 1996), and is often used for benchmarking inference schemes in high-dimensional settings. Following Drovandi et al. (2022), we set $x_{0}=(0,\ldots,0)^{\prime}$ and $\theta=(1,1,8,\sqrt{10})^{\prime}$. Two synthetic data sets, $\mathcal{D}_{1}$ (with $d=5$) and $\mathcal{D}_{2}$ (with $d=10$), were generated via recursive sampling of the Euler-Maruyama approximation of (29) over $[0,30]$ with $\Delta t=5\times 10^{-3}$. The resulting paths were thinned by a factor of 40 and corrupted with additive Gaussian noise, with zero mean and variance $R=5^{2}I_{d\times d}$, giving, for each of $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$, 30 observations with an inter-observation time of $0.2$ time units.
@@ 6 Discussion
Bayesian filtering for state-space models in the presence of unknown static parameters is a challenging problem. Sequential Monte Carlo algorithms must overcome particle degeneracy; one such scheme that deals with static parameters in a principled way is the SMC2 scheme of Chopin et al. (2013). Here, parameter particles are weighted by an estimate of the current observed-data likelihood and, subject to some degeneracy criterion, mutated via a (pseudo-marginal) Metropolis-Hastings kernel which has the marginal parameter posterior as its invariant distrubution. A particle filter over the latent state process is used to estimate the required observed-data likelihood contributions. However, use of the particle filter in this way is likely computationally prohibitive in high-dimensional observation settings, due to requiring that the number of state particles be scaled exponentially with dimension. On the other hand, the ensemble Kalman filter (EnKF, Evensen, 1994) shifts particles (known as ensemble members in the EnKF context) via suitable Gaussian approximations, thus avoiding resampling altogether. Consequently, there is now a vast literature demonstrating the empirical performance of the EnKF for high-dimensional state-space models (see e.g. Katzfuss et al., 2019, and the references therein).
@@
We have proposed to use the EnKF within the principled framework of SMC2 by weighting parameter particles by the EnKF likelihood and, when necessary, mutating these particles by using a Metropolis-Hastings step that again uses the EnKF likelihood. Extensions to the resulting nested EnKF (NEnKF) algorithm were considered. In particular, the computational cost of the mutation step can be lessened by screening proposed moves using a surrogate, computed locally using the most recent parameter samples and their likelihood evaluations. As with the simplest implementation of SMC2, NEnKF is plug-and-play, since only forward simulation of the latent state process and evaluation of the observation density are required. We found, across four applications of increasing observation dimension, that the NEnKF is able to provide accurate and computationally efficient inferences regarding the static parameters. Since the NEnKF combines a particle filter over parameters with an ensemble Kalman filter over states, we anticipate (and as evidenced by our numerical experiments) greatest utility of the proposed algorithm in settings where the number of observed dynamic components is large compared to the number of parameters to be estimated (e.g. $d_{\theta}<10$ and $d_{y}>10$). The simplest implementation of the NEnKF assumes a linear Gaussian observation model. In scenarios where this assumption is unreasonable, we have proposed to replace the EnKF with a particle filter that uses the EnKF to propagate state particles. Unlike related approaches (e.g. Papadakis et al., 2010), we also use the EnKF to give a linear Gaussian approximation of the transition density $p_{t}^{(\theta)}(x_{t}|x_{t-1})$, leading to a simple form of weight function that can be readily applied in settings where states are modelled at a finer frequency than observations (as is the case, for example, when working with discretised SDEs). The resulting Rao-Blackwellised particle filter can be used as the inner filter in SMC2, giving an inference scheme that is typically less accurate but more efficient than vanilla SMC2, and less efficient but more accurate than NEnKF.
@@
This work can be extended in a number of ways. For example, in our applications of the delayed acceptance kernel, the number of distinct parameter particles $M_{r}$ used to construct the $k$-$n$ surrogate was typically small (no more than a few thousand particles), justifying an $\mathcal{O}(M_{r})$ list look-up method for finding the $k$-nearest neighbours of each $\theta^{*}$. Using a variation of the KD-tree (e.g. Sherlock et al., 2017) requires $\mathcal{O}(d_{\theta}\log M_{r})$ operations and provides a cheaper alternative to the list look-up method, provided $d_{\theta}$ is moderate. It may also be possible to improve computational efficiency of the NEnKF by adaptively choosing between several delayed-acceptance kernels (with different $k$) and a standard M-H kernel, when the resample-move step is triggered. For example, in the early phase of the NEnKF scheme, calculating an observed-data likelihood estimator is likely to be relatively cheap, motivating use of a standard M-H kernel. Similar approaches have been considered in the context of IBIS (Fearnhead and Taylor, 2013) and SMC2 (Botha et al., 2023a). The efficiency of the resample-move step could be further improved by leveraging gradient-based proposals in the M-H step (e.g. Rosato et al., 2024). Finally, using the surrogate in both the weighting and resample-move steps (for a fixed value of $N$) could provide a fixed cost alternative to the NEnKF, to the detriment of inferential accuracy.